OpenStudy (anonymous):

If y=3/4+x^2 then dy/dx=

OpenStudy (anonymous):

\(2x\) ?

OpenStudy (anonymous):

unless you mean \(y=\frac{3}{4+x^2}\) in which case use \[\frac{d}{dx}[\frac{1}{f}]=-\frac{f'}{f^2}\] to get \[\frac{-6x}{(4+x^2)^2}\]

OpenStudy (anonymous):

OK